OpenStudy (aryana_maria2323):

For f(x) = ecos(x) use your graphing calculator to find the number of zeros for f '(x) on the closed interval [0, 2π]. a. 1 b. 2 c. 3 d. 4

OpenStudy (aryana_maria2323):

@Babynini

OpenStudy (babynini):

\[e^{cosx}\] is that what it is?

OpenStudy (aryana_maria2323):

Yes \[e ^{\cos \left( x \right)}\]

OpenStudy (babynini):

Ok so step one would be finding the derivative of that. Any ideas?

OpenStudy (aryana_maria2323):

I tried using the chain rule but still it didn't give me any of the answers. the i tried d/dx

OpenStudy (babynini):

\[y=e^{\cos(x)}\] so i would take ln of both sides..

OpenStudy (aryana_maria2323):

How would you do that?

OpenStudy (babynini):

\[\ln(y)=\ln(e^{\cos(x)})\] right?

OpenStudy (babynini):

and we just talked about how to solve the left side one. so what does the left side equal? :)

OpenStudy (aryana_maria2323):

Okay let me try that but it wants me to find the number of zeros for f(x) on the closed interval [0,2pie]

OpenStudy (babynini):

Well we need the derivative first though, because it wants the number of 0's for f'(x) not f(x)

OpenStudy (aryana_maria2323):

I got -y sin(x)

geerky42 (geerky42):

Ya so \(f'(x) = -\sin(x)\cdot e^{\cos(x)}\) Now use graphing calculator to graph \(f'(x)\) and count zeros.

OpenStudy (babynini):

Yeah! but y looks ugly there so plug in the original y y =e^{cos(x)}

OpenStudy (aryana_maria2323):

I think I did that wrong because I got 0 and that is not one of the answers

OpenStudy (aryana_maria2323):

?? @Babynini

OpenStudy (babynini):

Is meaning zeros meaning how many times it crosses the x axis?

OpenStudy (aryana_maria2323):

I believe so. The closed interval part [0,2pie] is throwing me off.

OpenStudy (babynini):

well here's the graph of it. and 2pi= 6.2ish

OpenStudy (babynini):

That spot where it crosses x after the 6 is exactly 2pi

OpenStudy (aryana_maria2323):

okay so 3 times?

OpenStudy (babynini):

yaya I think so:)

OpenStudy (aryana_maria2323):

Okay thank you. Do you think you have time for one more to help me with?

OpenStudy (babynini):

I can try.

OpenStudy (aryana_maria2323):

Graph the piece function using the values of a and b that you have found. You may graph by hand or use your calculator to graph and copy and paste into the document.

OpenStudy (aryana_maria2323):

 This is what the problems/ numbers are